I need help finding the difference quotient for the following function.
F(x)=5x-4
Find for these:
1) f(x+h)=
2) f(x)=
3)f(x+h)-f(x)=
4)(f(x+h)-f(x))/h=

Answer by Edwin McCravy(20081)

I need help finding the difference quotient for the following function.
f(x)=5x-4
Find for these:
1) f(x+h)=
2) f(x)=
3)f(x+h)-f(x)=
4)(f(x+h)-f(x))/h=
It's just a matter of substituting what's in the parentheses 
just after the function letter "f", in place of x in the original
equation.

    f(x)=5x-4

To find f(x+h) do this:

1. Write f(x)=5x-4 all spread out like this:

    f( x ) =  5x    -  4 

1. Take out all x's, leaving big spaces when you took 
them out, like this:

    f(   ) =  5     -  4

Now in those blank spaces where you took out the x's, 
put in (x+h) and you have this

    f(x+h) =  5(x+h)- 4

1)  f(x+h) =  5x+5h-4   <--simplified answer to 1)

2)  f(x) = 5x-4         <--answer to 2), that's given

    f(x+h)-f(x) = (5x+5h-4) - (5x-4)  <--be sure to use () here

    f(x+h)-f(x) = 5x+5h-4-5x+4

3)  f(x+h)-f(x) = 5h    <--simplified answer to 3)

    [f(x+h)-f(x)]/h = (5h)/h

4)  [f(x+h)-f(x)]/h = 5   <-- simplified answer to 4), which is 
                              just the number 5. 

Edwin